0 0
Add a comment Improve this question Transcribed image text
Answer #1

Let number of photon be N
Total energy = N* energy of 1 photon
          = N * h*c/lambda
    4.4*10^-3 = N * 6.626 *10^-34 *3*10^8/(525*10^-9)
N=1.162 *10^16   
Answer #2

Planck relationship:
E = h